OpenStudy (anonymous):

A bird flies 6 1/4 mi in 20 min. How long will it take the bird to fly 8 1/8 mi at the same speed? min.

OpenStudy (jhannybean):

Make a ratio.

OpenStudy (jhannybean):

Mi/min

OpenStudy (anonymous):

WEEEEEEEEEEEEEEEEEEEEEWWWWWWWWWWWWOOOOOOOOOOOOOOOOOOOOOOO Im Patrick

OpenStudy (anonymous):

WEEEWWWWWWWWWWOOOOOOOOOOWOWOWOWO

OpenStudy (anonymous):

Sorry I had to it's fun

OpenStudy (jhannybean):

\[\sf \frac{6~ \frac{1}{4}~ mi}{20~min} = \frac{8~\frac{1}{8}~mi}{x~ min}\]

OpenStudy (jhannybean):

Solve.

OpenStudy (anonymous):

10? I'm kinda dumb at ratios

OpenStudy (jhannybean):

First change \(\sf 6~ \frac{1}{4} = \frac{(6\cdot 4)+1}{4}\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

6.25

OpenStudy (jhannybean):

Do the same with \(\sf 8~\frac{1}{8} = \frac{(8\cdot 8)+1}{8}\)

OpenStudy (jhannybean):

Then cross multiply.

OpenStudy (anonymous):

8.125?

OpenStudy (anonymous):

I got 50.78125

OpenStudy (jhannybean):

Just keep it as a fraction.

OpenStudy (jhannybean):

Keep them both as fractions.

OpenStudy (anonymous):

I still don't get it ._. I'm sorry for being dumb

OpenStudy (anonymous):

@Bee_see @dtan5457

OpenStudy (bee_see):

I got 26. Do you have choices for the problem?